How many mg Ч Copper 2.15 x® are umol of copper а го ч
Answer #1

1 g = 10³ mg and 1 mol = 10^6 μmol

Molar mass of Copper (Cu) = 63.5 g/mol = (63.5 × 10³) mg/(10^6 μmol) = 0.0635 mg/μmol

Given moles of Cu = 2.15 × 10⁴ μmol

Mass of Cu = Moles × Molar mass = (2.15 × 10⁴ μmol) × (0.0635 mg/μmol) = 1365.25 mg

Hence Mass of Copper in its 2.15 × 10⁴ μmol is 1365.25 mg

If you find this solution helpful then please give a thumbsup ?. Thankyou.